# Efficient Unstructured Pruning of Mamba State-Space Models for Resource-Constrained Environments

### v1

@@ Abstract
State-space models (SSMs), particularly the Mamba architecture, have emerged as powerful alternatives to Transformers for sequence modeling, offering linear-time complexity and competitive performance across diverse tasks. However, their large parameter counts pose significant challenges for deployment in resource-constrained environments. We propose a novel unstructured pruning framework tailored for Mamba models that achieves up to 70% parameter reduction while retaining over 95% of the original performance. Our approach integrates three key innovations: (1) a gradient-aware magnitude pruning technique that combines weight magnitude and gradient information to identify less critical parameters, (2) an iterative pruning schedule that gradually increases sparsity to maintain model stability, and (3) a global pruning strategy that optimizes parameter allocation across the entire model. Through extensive experiments on WikiText-103, Long Range Arena, and ETT time-series benchmarks, we demonstrate significant efficiency gains with minimal performance degradation. Our analysis of pruning effects on Mamba’s components reveals critical insights into the architecture’s redundancy and robustness, enabling practical deployment in resource-constrained settings while broadening Mamba’s applicability.
@@ 1 Introduction
Sequence modeling has been revolutionized by attention-based Transformers [1, 2, 3], yet these architectures struggle with quadratic computational complexity [4], limiting their use in long-context tasks and resource-constrained environments. State-space models (SSMs) [5, 6, 7] offer a promising alternative with linear-time complexity while effectively modeling long-range dependencies.
The Mamba architecture [8] distinguishes itself through its selective mechanism that dynamically controls information flow based on input data. This has led to state-of-the-art performance across language modeling [9], time-series forecasting [10], audio processing [11], and long-context understanding [12]. Mamba’s recurrent formulation avoids memory-intensive attention matrices, enabling efficient computation through convolution-like operations. This favorable scaling has spurred extensions to vision [13], multimodal processing [14], and genomics [15]. Despite these advances, deploying Mamba models in resource-constrained environments remains challenging due to their millions of parameters [16]. Neural network pruning offers a potential solution, but techniques developed for CNNs [17, 18] or Transformers [19, 20] don’t directly transfer to Mamba’s unique recurrent structure and state-space dynamics [21, 22].
We introduce a systematic unstructured pruning framework tailored to Mamba’s architecture, enabling deployment in resource-constrained settings like edge computing and mobile devices. Our approach combines three innovations: (1) gradient-aware magnitude pruning that identifies less important parameters while preserving model expressiveness; (2) an iterative pruning schedule with cubic progression that gradually increases sparsity; and (3) a global pruning strategy that optimizes parameter allocation across the entire model. Experiments on WikiText-103 [9], Long Range Arena [12], and ETT [10] demonstrate up to 70% parameter reduction with over 95% performance retention. Our contributions include:
@@ 2 Related Work
Our work builds on advancements in state-space models (SSMs) and neural network pruning, tailoring these techniques to the unique properties of the Mamba architecture. Below, we summarize the most relevant literature, with a broader review of sequence modeling architectures provided in Appendix D.
Neural Network Pruning. Pruning reduces model size by removing redundant parameters, with early work using second-order derivatives [23, 24] and later approaches focusing on magnitude-based pruning [25, 26]. The lottery ticket hypothesis [27] showed that sparse subnetworks can match dense model performance. Pruning has been applied to CNNs [17, 18] and Transformers [19, 20], but these methods do not account for the recurrent dynamics of SSMs [21]. Gradient-based pruning [28, 29, 30], which considers both weight magnitude and gradient information, shows promise but has not been extensively explored for SSMs.
@@ 3 Methodology
To enable efficient deployment of Mamba state-space models in resource-constrained environments, we propose a comprehensive unstructured pruning framework tailored to their unique architecture. Our approach addresses the challenges of preserving Mamba’s selective mechanism and stable recurrent dynamics while significantly reducing parameter counts. Figure 1 provides an overview of our approach.
Figure 1: Overview of our unstructured pruning framework for Mamba models. The method combines (1) gradient-aware magnitude pruning, which uses weight magnitude and gradient information to compute importance scores, (2) an iterative pruning schedule with cubic progression to ensure stability, and (3) global pruning to optimize parameter allocation across all layers. The three plots show: (left) the gradient-aware importance distribution across parameters, (middle) the cubic sparsity progression over training iterations, and (right) the model performance vs. sparsity trade-off compared to baseline methods.
@@ 3.1.1 Gradient-Aware Magnitude Pruning
The core of our pruning strategy is a gradient-aware magnitude pruning technique that identifies parameters with minimal impact on model performance. While this approach builds upon insights from previous gradient-based pruning methods like SNIP [28] and magnitude pruning [25], our formulation and application are specifically tailored to Mamba’s unique architecture. Unlike traditional magnitude-based pruning, which solely considers weight magnitude, our method incorporates gradient information to assess a parameter’s contribution to the loss function, ensuring that critical parameters are preserved. For each parameter $w_{ij}$ in the Mamba model, we compute an importance score $S(w_{ij})$ defined as:
@@
Here, $|w_{ij}|$ is the absolute weight magnitude, $\frac{\partial\mathcal{L}}{\partial w_{ij}}$ is the gradient of the loss $\mathcal{L}$ with respect to $w_{ij}$, and $\alpha$ is a tunable hyperparameter that balances the influence of magnitude and gradient. A value of $\alpha=0$ reduces to pure magnitude pruning, while $\alpha>0$ emphasizes parameters with significant impact on the loss. Through extensive hyperparameter sweeps (detailed in Appendix B), we find that $\alpha\approx 1.0$ provides a robust default across tasks, as it equally weighs magnitude and gradient contributions, though task-specific tuning can yield further improvements (e.g., $\alpha\approx 0.8$ for time-series forecasting).
@@ 3.1.2 Iterative Pruning Schedule
Rather than pruning all parameters at once, we employ an iterative schedule that gradually increases sparsity over training. Building upon the cubic pruning schedule proposed by Zhu et al. [26], we adapt this approach specifically for Mamba’s recurrent dynamics and selective attention mechanism. This gradual pruning allows the remaining parameters to compensate for the pruned ones, leading to better recovery of performance. Given an initial sparsity level $s_{0}$ (usually 0), a final target sparsity level $s_{f}$, and pruning starting at iteration $t_{0}$ and continuing until total training iteration $T$, the sparsity at iteration $t$ follows a cubic progression:
$s_{t}=s_{f}+(s_{0}-s_{f})\cdot\left(1-\frac{t-t_{0}}{T-t_{0}}\right)^{3}\quad\text{for}\quad t\in[t_{0},T]$ (2)
Here, $t_{0}$ is the iteration to start pruning (typically after 25% of training), $T$ is the total training iterations, and the cubic term ensures a gradual initial phase followed by accelerated pruning. This schedule starts slowly, allowing the model to converge toward important parameter configurations, and then accelerates to reach the target sparsity. Our empirical findings (detailed in Appendix C) demonstrate that this cubic schedule is particularly effective for Mamba models compared to linear or exponential alternatives. We hypothesize this is due to Mamba’s recurrent nature, where parameter interactions are more complex than in feed-forward architectures, requiring a more gradual initial pruning phase to maintain stability of the state dynamics while allowing sufficient adaptation time before reaching high sparsity.
@@ 3.1.3 Global Pruning Strategy
Traditional pruning methods often apply layer-wise thresholds, which can lead to suboptimal parameter allocation by treating each layer independently [25]. In contrast, our global pruning strategy computes a single importance threshold across all parameters in the Mamba model, allowing for flexible and efficient distribution of sparsity. After computing importance scores $S(w_{ij})$ for all parameters, we sort them globally and mask the lowest-scoring parameters to achieve the target sparsity level. This global approach is particularly effective for Mamba, as its architecture exhibits varying parameter importance across layers and components (e.g., state-space vs. linear projections). For example, earlier layers, which capture foundational features, often retain more parameters than later layers, as shown in Appendix C.2. Global pruning outperforms layer-wise pruning by up to 0.5 perplexity points on language modeling tasks (see Appendix C), as it optimizes the overall model capacity rather than enforcing uniform sparsity per layer.
@@ 3.1.4 Eigenvalue Stability Preservation
where $\lambda_{i,j}$ is the eigenvalue of the transition matrix for state dimension $i$ at position $j$, and $\epsilon$ is a small positive value (typically 0.01) providing a safety margin. Parameters that minimize $S_{\text{stab}}$ are preferentially retained to maintain stability. In practice, this is implemented as a corrective mechanism that adjusts the pruning mask post-hoc if stability violations are detected, preventing the removal of parameters critical for maintaining eigenvalue bounds (see Algorithm 1 in Appendix E for details). This stability-aware pruning ensures that the model’s recurrent dynamics remain well-behaved even at high sparsity levels.
@@
Our pruning framework is implemented in PyTorch, wrapping the Mamba model with a pruning mask that enables sparse matrix operations during training and inference. The importance scores are computed using gradients from a single forward-backward pass per pruning step, minimizing computational overhead. We use the AdamW optimizer [31] for fine-tuning after each pruning step, with a learning rate schedule that decreases linearly from $10^{-4}$ to $10^{-6}$. The hyperparameter $\alpha$ is tuned via a grid search over $[0,0.5,1.0,2.0]$, with task-specific sweeps detailed in Appendix B. Sparse operations leverage PyTorch’s sparse tensor support, reducing memory usage by up to 54% at 50% sparsity (see Appendix C.2). The framework is compatible with various Mamba variants (e.g., Vision Mamba [13], Hyena [32]), demonstrating its generality across state-space architectures.
@@ 4 Results
We evaluate our unstructured pruning framework on Mamba models across diverse tasks, including language modeling, long-range understanding, and time-series forecasting, using benchmark datasets such as WikiText-103 [9], Long Range Arena [12], and ETT [10]. Our experiments demonstrate that the proposed approach achieves up to 70% parameter reduction with minimal performance degradation, significantly outperforming baseline pruning methods. We also analyze computational efficiency and robustness, highlighting the practical benefits for resource-constrained deployment. A component-wise analysis reveals critical insights into Mamba’s pruning characteristics. Extended results, including fine-grained ablations and cross-dataset performance, are provided in Appendices C.2 and C.
@@ 4.1 Language Modeling Performance
We assess our pruning framework on language modeling using WikiText-103 and PG-19 datasets, comparing pruned Mamba models to dense baselines and conventional magnitude-based pruning [25]. Table 1 summarizes the results for Mamba-Small (130M parameters) and Mamba-Base (370M parameters).
Table 1: Language modeling perplexity (lower is better) and inference time on WikiText-103 and PG-19. Parameter counts are in millions. Inference time is measured in milliseconds per token on a single NVIDIA A100 GPU.

### v2

@@ Abstract
As AI deployment shifts to edge devices, efficient sequence modeling becomes critical. State-space models (SSMs), particularly Mamba, rival Transformers with linear-time complexity and strong performance across tasks, yet their large parameter counts hinder resource-constrained use. We propose a novel unstructured pruning framework tailored for Mamba, achieving up to 70% parameter reduction with only 3–9% performance loss. Unlike Transformer-focused pruning, our approach leverages Mamba’s recurrent dynamics through: (1) pruning based on weight and gradient importance to preserve critical parameters, (2) a gradual pruning process to ensure model stability, and (3) a global strategy optimizing parameter allocation across the model. Extensive experiments on WikiText-103, Long Range Arena, and ETT benchmarks show significant efficiency gains, with 1.77× faster inference and 46% less memory. Our component analysis reveals Mamba’s robustness, enabling practical deployment while requiring careful use to avoid biases in sensitive applications.
Efficient Unstructured Pruning of Mamba State-Space Models for Resource-Constrained Environments
Ibne Farabi Shihab1*, Sanjeda Akter1*, Anuj Sharma2 1Department of Computer Science, Iowa State University, Ames, IA, USA, 2Department of Civil, Construction and Environmental Engineering, Iowa State University, Ames, IA, USA *Equal contribution. Correspondence: ishihab@iastate.edu
@@ 1 Introduction
Sequence modeling has been revolutionized by attention-based Transformers Vaswani et al. (2017); Devlin et al. (2018); Brown et al. (2020), yet these architectures struggle with quadratic computational complexity Tay et al. (2022a), limiting their use in long-context tasks and resource-constrained environments. State-space models (SSMs) Gu et al. (2020a, 2021); Gupta et al. (2022) offer a promising alternative with linear-time complexity while effectively modeling long-range dependencies.
The Mamba architecture Gu and Dao (2023) distinguishes itself through its selective mechanism that dynamically controls information flow based on input data. This has led to state-of-the-art performance across language modeling Merity et al. (2016), time-series forecasting Zhou et al. (2021), audio processing Goel et al. (2022), and long-context understanding Tay et al. (2020). Mamba’s recurrent formulation avoids memory-intensive attention matrices, enabling efficient computation through convolution-like operations. This favorable scaling has spurred extensions to vision Zhu et al. (2024), multimodal processing Qiao et al. (2024), and genomics Nguyen et al. (2023).
Despite these advances, deploying Mamba models in resource-constrained environments remains challenging due to their millions of parameters Deng et al. (2020). Neural network pruning offers a potential solution, but techniques developed for CNNs Li et al. (2016); He et al. (2018) or Transformers Michel et al. (2019); Voita et al. (2019) don’t directly transfer to Mamba’s unique recurrent structure and state-space dynamics Liu et al. (2021); Bellec et al. (2018).
We introduce a systematic unstructured pruning framework tailored to Mamba’s architecture, enabling deployment in resource-constrained settings like edge computing and mobile devices. Our approach combines three innovations: (1) gradient-aware magnitude pruning that identifies less important parameters while preserving model expressiveness; (2) an iterative pruning process to ensure model stability during sparsity increases; and (3) a global pruning strategy that optimizes parameter allocation across the entire model. Experiments on WikiText-103 Merity et al. (2016), Long Range Arena Tay et al. (2020), and ETT Zhou et al. (2021) demonstrate up to 70% parameter reduction with only 3-9% performance degradation.
Our contributions include:
@@
Figure 1: Detailed workflow illustrating our unstructured pruning framework for Mamba models. It includes (1) gradient-aware magnitude pruning with importance scores that balance weight magnitude and gradient information, (2) an iterative cubic pruning schedule that gradually increases sparsity, and (3) stability-preserving mechanisms that maintain eigenvalue bounds. This process optimizes global parameter allocation to achieve 50–70% parameter reduction with minimal performance loss.
@@ 2 Related Work
Our work builds on advancements in state-space models (SSMs) and neural network pruning, tailoring these techniques to the unique properties of the Mamba architecture. Below, we summarize the most relevant literature, with a broader review of sequence modeling architectures provided in Appendix E.
Neural Network Pruning. Pruning reduces model size by removing redundant parameters, with early work using second-order derivatives LeCun et al. (1990); Hassibi et al. (1993) and later approaches focusing on magnitude-based pruning Han et al. (2015); Zhu and Gupta (2017). The lottery ticket hypothesis Frankle and Carbin (2018) showed that sparse subnetworks can match dense model performance. Pruning has been applied to CNNs Li et al. (2016); He et al. (2018) and Transformers Michel et al. (2019); Voita et al. (2019), but these methods do not account for the recurrent dynamics of SSMs Liu et al. (2021). Recent advances in large language model pruning include simple weight-magnitude methods like Wanda Sun et al. (2024), which are effective for Transformers but don’t address the unique stability requirements of recurrent state-space models. Gradient-based pruning Lee et al. (2018); Wang et al. (2020a); Molchanov et al. (2019), which considers both weight magnitude and gradient information, shows promise but has not been extensively explored for SSMs.
@@ 3 Methodology
To enable efficient deployment of Mamba state-space models in resource-constrained environments, we propose a comprehensive unstructured pruning framework tailored to their unique architecture, figure 1 illustrates the framework workflow. Our approach addresses the challenges of preserving Mamba’s selective mechanism and stable recurrent dynamics while significantly reducing parameter counts. Figure 2 provides an overview of our approach.
Figure 2: Our unstructured pruning framework for Mamba models combines (1) gradient-aware magnitude pruning, (2) an iterative cubic pruning schedule, and (3) global parameter allocation. The diagram illustrates importance distribution, sparsity progression, and performance trade-offs.
@@ 3.1.1 Gradient-Aware Magnitude Pruning
The core of our pruning strategy is a gradient-aware magnitude pruning technique that identifies parameters with minimal impact on model performance. While this approach builds upon insights from previous gradient-based pruning methods like SNIP Lee et al. (2018) and magnitude pruning Han et al. (2015), our formulation and application are specifically tailored to Mamba’s unique architecture. Unlike traditional magnitude-based pruning, which solely considers weight magnitude, our method incorporates gradient information to assess a parameter’s contribution to the loss function, ensuring that critical parameters are preserved. For each parameter $w_{ij}$ in the Mamba model, we compute an importance score $S(w_{ij})$ defined as:
@@
Here, $|w_{ij}|$ is the absolute weight magnitude, $\frac{\partial\mathcal{L}}{\partial w_{ij}}$ is the gradient of the loss $\mathcal{L}$ with respect to $w_{ij}$, and $\alpha$ is a tunable hyperparameter that balances the influence of magnitude and gradient. A value of $\alpha=0$ reduces to pure magnitude pruning, while $\alpha>0$ emphasizes parameters with significant impact on the loss. Through extensive hyperparameter sweeps (detailed in Appendix C), we find that $\alpha\approx 1.0$ provides a robust default across tasks, as it equally weighs magnitude and gradient contributions, though task-specific tuning can yield further improvements (e.g., $\alpha\approx 0.8$ for time-series forecasting). The impact of different $\alpha$ values on performance across tasks is presented in Appendix C.
@@ 3.1.2 Iterative Pruning Schedule
Rather than pruning all parameters at once, we employ an iterative schedule that gradually increases sparsity over training. Building upon the cubic pruning schedule proposed by Zhu et al. Zhu and Gupta (2017), we adapt this approach specifically for Mamba’s recurrent dynamics and selective attention mechanism. This gradual pruning allows the remaining parameters to compensate for the pruned ones, leading to better recovery of performance. Given an initial sparsity level $s_{0}$ (usually 0), a final target sparsity level $s_{f}$, and pruning starting at iteration $t_{0}$ and continuing until total training iteration $T$, the sparsity at iteration $t$ follows a cubic progression:
$s_{t}$ $=s_{f}+(s_{0}-s_{f})\cdot\ \left(1-\frac{t-t_{0}}{T-t_{0}}\right)^{3},$ (2)
$\quad\text{for }t\in[t_{0},T]$
Here, $t_{0}$ is the iteration to start pruning (typically after 25% of training), $T$ is the total training iterations, and the cubic term ensures a gradual initial phase followed by accelerated pruning. This schedule starts slowly, allowing the model to converge toward important parameter configurations, and then accelerates to reach the target sparsity. Our empirical findings (detailed in Appendix C) demonstrate that this cubic schedule is particularly effective for Mamba models compared to linear or exponential alternatives. A comparative analysis of different pruning schedules can be found in Appendix C.
@@ 3.1.3 Global Pruning Strategy
Traditional pruning methods often apply layer-wise thresholds, which can lead to suboptimal parameter allocation by treating each layer independently Han et al. (2015). In contrast, our global pruning strategy computes a single importance threshold across all parameters in the Mamba model, allowing for flexible and efficient distribution of sparsity. After computing importance scores $S(w_{ij})$ for all parameters, we sort them globally and mask the lowest-scoring parameters to achieve the target sparsity level. This global approach is particularly effective for Mamba, as its architecture exhibits varying parameter importance across layers and components (e.g., state-space vs. linear projections). For example, earlier layers, which capture foundational features, often retain more parameters than later layers, as shown in Appendix C.1. Global pruning outperforms layer-wise pruning by up to 0.5 perplexity points on language modeling tasks (see Appendix C), as it optimizes the overall model capacity rather than enforcing uniform sparsity per layer.
@@ 3.1.4 Eigenvalue Stability Preservation
where $\lambda_{i,j}$ is the eigenvalue of the transition matrix for state dimension $i$ at position $j$, and $\epsilon$ is a small positive value (typically 0.01) providing a safety margin. Parameters that minimize $S_{\text{stab}}$ are preferentially retained to maintain stability. In practice, this is implemented as a corrective mechanism that adjusts the pruning mask post-hoc if stability violations are detected, preventing the removal of parameters critical for maintaining eigenvalue bounds (see Algorithm 1 in Appendix D.6 for details). This stability-aware pruning ensures that the model’s recurrent dynamics remain well-behaved even at high sparsity levels.
@@
Our pruning framework is implemented in PyTorch, wrapping the Mamba model with a pruning mask that enables sparse matrix operations during training and inference. The importance scores are computed using gradients from a single forward-backward pass per pruning step, minimizing computational overhead. We use the AdamW optimizer Loshchilov and Hutter (2017) for fine-tuning after each pruning step, with a learning rate schedule that decreases linearly from $10^{-4}$ to $10^{-6}$. The hyperparameter $\alpha$ is tuned via a grid search over $[0,0.5,1.0,2.0]$, with task-specific sweeps detailed in Appendix C. Sparse operations leverage PyTorch’s sparse tensor support, reducing memory usage by up to 54% at 50% sparsity (see Appendix C.1). The framework is compatible with various Mamba variants (e.g., Vision Mamba Zhu et al. (2024), Hyena Poli et al. (2023)), demonstrating its generality across state-space architectures.
3.2 Theoretical Foundations
Our pruning approach is grounded in theoretical insights about Mamba’s architecture and the unique challenges of pruning recurrent state-space models. These insights inform both the design of our framework and explain its effectiveness.
3.2.1 Parameter Importance Distribution
The distribution of parameter importance in Mamba follows a power law, with a small fraction of parameters contributing disproportionately to model performance. Our analysis shows that approximately 20% of parameters account for 80% of the total importance score, creating a natural opportunity for high-sparsity pruning. This power law distribution arises from Mamba’s selective mechanism, which creates context-dependent parameter activation patterns where different inputs activate distinct parameter subsets.
Unlike Transformers, where attention weights tend to be distributed more uniformly, Mamba’s recurrent structure leads to more concentrated parameter importance as many parameters serve similar roles. The state-space parameters (A, B, C, D matrices) exhibit higher importance and enable targeted pruning. The effective rank of activation matrices is lower than in Transformers, indicating greater redundancy exploitable by pruning.
3.2.2 Eigenvalue Stability Theory
For recurrent models like Mamba, maintaining eigenvalue stability during pruning is essential. Using matrix perturbation theory, we can quantify the maximal eigenvalue shift when pruning a state transition matrix $A$ to its pruned counterpart $\tilde{A}$:
where $s$ is sparsity, $\|\bar{A}\|_{F}$ is the Frobenius norm, and $C$ depends on matrix structure. At 50% sparsity, our method ensures the maximum shift is $\leq 0.05$, preserving the stability necessary for effective sequence modeling.
3.3 Mamba-Specific Adaptations
While the pruning techniques described above build on established principles, their effectiveness in our framework comes from specific adaptations to Mamba’s unique architecture, directly addressing its recurrent dynamics, selective gating, and stability requirements. These adaptations substantiate our claim that this is a Mamba-tailored pruning framework.
3.3.1 Importance Scoring for Recurrent Dynamics
Standard gradient-based importance scores like SNIP are typically computed in a single forward-backward pass. For a recurrent model like Mamba, this fails to capture a parameter’s influence across an entire sequence. We adapt the gradient calculation to account for this temporal dependency. The gradient $\frac{\partial\mathcal{L}}{\partial w_{ij}}$ used in Equation 1 is accumulated over multiple time steps of the recurrent computation, providing a more holistic measure of a parameter’s contribution to the sequence-level loss:
$S_{\text{SSM}}(w_{ij})=|w_{ij}|\cdot\left|\sum_{t=1}^{T}\frac{\partial\mathcal{L}}{\partial w_{ij}}_{(t)}\right|^{\alpha}$ (5)
where the gradient is accumulated across $T$ time steps, capturing how parameters influence the model across the entire sequence rather than at isolated points.
3.3.2 Preserving the Selective Mechanism
Mamba’s performance relies heavily on its selective gating mechanism, which is data-dependent. We found that conventional importance scores systematically underestimated the importance of gating parameters that enable this selectivity. To address this, we introduce a correction factor to the importance score for gating parameters, which scales their importance based on the diversity of their activation patterns across different inputs. This ensures that parameters critical for dynamic, input-dependent information routing are preserved.
3.3.3 Stability-Aware Scheduling
The cubic pruning schedule is adapted to prevent the destabilization of Mamba’s recurrent dynamics. After each pruning step, the fine-tuning process incorporates a stability-focused objective that explicitly penalizes eigenvalues of the state transition matrix that drift outside the unit circle:
$\mathcal{L}_{\text{stable}}=\mathcal{L}_{\text{task}}+\lambda\cdot\sum_{i}\max(0,|\lambda_{i}|-(1-\epsilon))^{2}$ (6)
This dual-objective fine-tuning allows the model to recover task performance while ensuring its recurrent states remain stable, a critical consideration absent in frameworks designed for non-recurrent models. Further theoretical details on these adaptations are available in Appendix D.8.
@@ 4 Results
We evaluate our unstructured pruning framework on Mamba models across diverse tasks, including language modeling, long-range understanding, and time-series forecasting, using benchmark datasets such as WikiText-103 Merity et al. (2016), Long Range Arena Tay et al. (2020), and ETT Zhou et al. (2021). Our experiments demonstrate that the proposed approach achieves up to 70% parameter reduction with minimal performance degradation, significantly outperforming baseline pruning methods. We also analyze computational efficiency and robustness, highlighting the practical benefits for resource-constrained deployment. A component-wise analysis reveals critical insights into Mamba’s pruning characteristics. Extended results, including fine-grained ablations and cross-dataset performance, are provided in Appendix C.
@@ 4.1 Language Modeling Performance
We assess our pruning framework on language modeling using WikiText-103 and PG-19 datasets, comparing pruned Mamba models to dense baselines and conventional magnitude-based pruning Han et al. (2015). Table 1 summarizes the results for Mamba-Small (130M parameters) and Mamba-Base (370M parameters).
Table 1: Language modeling perplexity and inference time on WikiText-103 and PG-19 for models with varying parameter counts and sparsity levels.